# PromptVFX: Text-Driven Fields for Open-World 3D Gaussian Animation

### v1

@@ Abstract
Visual effects (VFX) are key to immersion in modern films, games, and AR/VR. Creating 3D effects requires specialized expertise and training in 3D animation software and can be time consuming. Generative solutions typically rely on computationally intense methods such as diffusion models which can be slow at 4D inference. We reformulate 3D animation as a field prediction task and introduce a text-driven framework that infers a time-varying 4D flow field acting on 3D Gaussians. By leveraging large language models (LLMs) and vision-language models (VLMs) for function generation, our approach interprets arbitrary prompts (e.g., “make the vase glow orange, then explode”) and instantly updates color, opacity, and positions of 3D Gaussians in real time. This design avoids overheads such as mesh extraction, manual or physics-based simulations and allows both novice and expert users to animate volumetric scenes with minimal effort on a consumer device even in a web browser  [^1] Web Demo: http://promptvfx.duckdns.org. Experimental results show that simple textual instructions suffice to generate compelling time-varying VFX, reducing the manual effort typically required for rigging or advanced modeling. We thus present a fast and accessible pathway to language-driven 3D content creation that can pave the way to democratize VFX further. Code available at 3dwe-ai.github.io/promptvfx.
1 Introduction
PromptVFX. In this paper, we propose PromptVFX, a text-driven framework for 3D animation that provides these features. Centered on Gaussian Splatting [10], our system translates high-level textual instructions into time-varying transformations on a set of Gaussians without generating new geometry or launching physics simulators. A time-varying field is applied to a Gaussian splatting scene ultimately changing centre positions, colour and opacity of individual Gaussians. When instructed to “raise the vase by two meters,” the system directly modifies the Gaussian centre positions over time. This simple yet generic idea naturally supports real-time rendering such that the effect of a user prompt can be visualized within seconds rather than minutes or hours.
@@ 3 Methodology
Our goal is to enable real-time, text-driven animation of 3D scenes without relying on diffusion-based optimization or physics simulators. In this section, we describe our approach in four parts: (1) representation of 3D scenes as Gaussian splats, (2) formulation of continuous time-dependent animation fields, (3) LLM-based translation of open-domain text into parametric fields, and (4) pipeline details ensuring efficient, user-friendly performance. Figure 1 provides an overview of our zero-shot pipeline.
@@ 3.2 Time-Dependent Fields for Animation
where $i$ indicates a particular Gaussian, and $t\in[0,T],T\in\mathbb{R}$ spans the animation’s duration. These fields allow structured transformations such as translation, rotation, color transitions, or opacity changes over time.
@@ 3.3.3 Animation Hypotheses Generation & Scoring
Each variation applies slightly different motion curves or color transitions. We then render each candidate animation (producing, for instance, a short sequence of frames) and feed these frames, along with the original text prompt, to a VLM that scores how well each animation matches the intended description on a 00100100100100 scale. We select the highest-scoring candidate as our base animation. We also include two optional refinement processes to improve the output.
@@ 3.3.5 Conversational Refinement
Table 2: CLIP and VQAScore scores for each method on different prompts.
Ours Gaussians2Life [5] AutoVFX [8]
Prompt CLIP [58] VQAScore [59] CLIP [58] VQAScore [59] CLIP [58] VQAScore [59]
Turn the vase into lava. 0.1712 0.7150 0.2058 0.1962 0.1587 0.4437
The bear expands and contracts like a lung. 0.2044 0.2833 0.2381 0.2583 0.2393 0.2832
The bulldozer accelerates forward. 0.2272 0.6845 0.3084 0.6813 0.1717 0.4512
@@ 4.3 Quantitative Evaluation
[…]
Results. Table 2 shows that PromptVFX achieves the highest VQAScore [59] across all evaluated prompts, demonstrating better text-video alignment from a temporal and semantic perspective. While Gaussians2Life [5] ranks higher in CLIP [58] for some cases due to preserving frame-level appearance, it does not always reflect coherent or realistic motion.
@@ 4.4.4 Effect of VLM Choice on Animation Quality
We compare different vision-language models used in our pipeline. Figure 7 illustrates that GPT-4o-mini produces a fluent bulldozer acceleration, while LLaVA [63, 64] struggles with maintaining object consistency and motion realism. This suggests that model selection significantly affects animation quality and highlights the need for well-designed system prompts tailored to the VLM’s capabilities.

### v2

@@ Abstract
Visual effects (VFX) are key to immersion in modern films, games, and AR/VR. Creating 3D effects requires specialized expertise and training in 3D animation software and can be time consuming. Generative solutions typically rely on computationally intense methods such as diffusion models which can be slow at 4D inference. We reformulate 3D animation as a field prediction task and introduce a text-driven framework that infers a time-varying 4D flow field acting on 3D Gaussians. By leveraging large language models (LLMs) and vision-language models (VLMs) for function generation, our approach interprets arbitrary prompts (e.g., “make the vase glow orange, then explode”) and instantly updates color, opacity, and positions of 3D Gaussians in real time. This design avoids overheads such as mesh extraction, manual or physics-based simulations and allows both novice and expert users to animate volumetric scenes with minimal effort on a consumer device even in a web browser  [^1] Web Demo: http://promptvfx.duckdns.org. Experimental results show that simple textual instructions suffice to generate compelling time-varying VFX, reducing the manual effort typically required for rigging or advanced modeling. We thus present a fast and accessible pathway to language-driven 3D content creation that can pave the way to democratize VFX further. Code available at https://obsphera.github.io/promptvfx/.
11footnotetext: Authors contributed equally to this work.
@@
1 Introduction
In this paper, we propose PromptVFX, a text-driven framework for 3D animation that provides these features. Centered on Gaussian Splatting [22], our system translates high-level textual instructions into time-varying transformations on a set of Gaussians without generating new geometry or launching physics simulators. A time-varying field is applied to a Gaussian splatting scene ultimately changing centre positions, colour and opacity of individual Gaussians. When instructed to “raise the vase by two meters,” the system directly modifies the Gaussian centre positions over time. This simple yet generic idea naturally supports real-time rendering such that the effect of a user prompt can be visualized within seconds rather than minutes or hours.
@@ 3 Methodology
Our goal is to enable real-time, text-driven animation of 3D scenes without relying on diffusion-based optimization or physics simulators. In this section, we describe our approach in four parts:
  1. (1)
    representation of 3D scenes as Gaussian splats,
  2. (2)
    formulation of continuous time-dependent animation fields,
  3. (3)
    LLM-based translation of open-domain text into parametric fields, and
  4. (4)
    pipeline details ensuring efficient, user-friendly performance.
Figure 1 provides an overview of our zero-shot pipeline.
@@ 3.2 Time-Dependent Fields for Animation
where $i$ indicates a particular Gaussian, and $t\in[0,T],T\in\mathbb{R}$ spans the animation’s duration. These fields allow structured transformations such as translation, color transitions, or opacity changes over time.
@@ 3.3.3 Animation Hypotheses Generation & Scoring
Each variation applies slightly different motion curves or color transitions. We then render each candidate animation (producing, for instance, a short sequence of frames) and feed these frames, along with the original text prompt, to a VLM that scores how well each animation matches the intended description on a 0100100 scale. We select the highest-scoring candidate as our base animation. We also include two optional refinement processes to improve the output.
@@
Table 2: Comparison of methods across CLIP similarity, VQAScore, and User Study ratings. User Study includes Text Alignment and Animation Quality scores (1–5 Likert scale, normalized). Bold indicates best per prompt and metric.
Prompt Method CLIP [40] VQAScore [28] User Study
Text Alignment Animation Quality
Turn the vase into lava. Gaussians2Life [50] 0.206 0.196 0.043 0.235
AutoVFX [18] 0.159 0.444 0.615 0.485
Ours 0.171 0.715 0.750 0.593
The bear expands and contracts like a lung. Gaussians2Life [50] 0.238 0.258 0.235 0.285
AutoVFX [18] 0.239 0.283 0.073 0.315
Ours 0.204 0.283 0.693 0.465
The bulldozer accelerates forward. Gaussians2Life [50] 0.308 0.681 0.278 0.235
AutoVFX [18] 0.172 0.451 0.035 0.058
Ours 0.227 0.685 0.893 0.685
@@ 4.3 Quantitative Evaluation
Evaluating 3D animation quality is difficult due to the absence of ground-truth sequences. Following prior work [8, 18, 50], we report CLIP similarity [40] as a frame-level proxy for text-to-image alignment. However, CLIP does not capture motion or temporal coherence. We therefore also report VQAScore, a video-level metric shown to better align with human judgment [28]. It estimates the probability that a model answers “Yes” to the prompt: Does this video align with the described animation: “{prompt}”?, denoted as $\mathbb{P}(\text{``Yes''}\mid\text{video, prompt})$.
Table 2 shows that PromptVFX achieves the highest VQAScore [28] across all evaluated prompts, demonstrating better text-video alignment from a temporal and semantic perspective. While Gaussians2Life [50] ranks higher in CLIP [40] for some cases due to preserving frame-level appearance, it does not always reflect coherent or realistic motion.
To further evaluate perceptual quality, we conducted a user study with 35 participants, following the protocol from AutoVFX [18]. Participants rated the generated videos on two criteria: Text Alignment and Animation Quality, using a 1-5 Likert scale (normalized to [0, 1] using min-max normalization). As shown in Table 2, PromptVFX outperforms all baselines across both criteria and all prompts, demonstrating its effectiveness in producing coherent and semantically grounded 3D animations.
@@ 4.4.3 Effect of VLM Choice on Animation Quality
We compare different vision-language models used in our pipeline. Figure 6 illustrates that GPT-4o-mini produces a fluent bulldozer acceleration, while LLaVA [31, 30] struggles with maintaining object consistency and motion realism.
This suggests that model selection significantly affects animation quality and highlights the need for well-designed system prompts tailored to the VLM’s capabilities.
4.4.4 Effect of Hypothesis Sampling